OpenStudy (anonymous):

A water pump takes 70 minutes to fill a tank. A larger pump can fill the same tank in 30 minutes. How long will it take to fill the tank if both pumps are used? a. 15 minutes b. 21 minutes c. 40 minutes d. 50 minutes

OpenStudy (anonymous):

I believe it would be 21 minutes.

OpenStudy (anonymous):

That's what I thought too.

OpenStudy (***[isuru]***):

the that will take by both pumps can be written as.. \[\frac{ 1 }{ t1} +\frac{ 1 }{ t2 } = \frac{ 1 }{ t3}\] where t1 = time by small pump t2 = time by large pump t3 = time by both pumps now substitute the data... \[\frac{ 1 }{ 70} +\frac{ 1 }{ 30} = \frac{ 1 }{ t3}\] ( 30+ 70) /2100= 1/t3 t3 = 2100/ 100 = 21 minutes hope this will help ya!!

OpenStudy (anonymous):

Lol very nice, I like how you described it!

OpenStudy (***[isuru]***):

we all make mistakes bro.. it should be corrected as * time that will take by both ....